# Inference in pseudo-observation-based regression using (biased) covariance estimation and naive bootstrapping

### v1

@@ Abstract
We demonstrate that the usual Huber-White estimator is not consistent for the limiting covariance of parameter estimates in pseudo-observation regression approaches. By confirming that a plug-in estimator can be used instead, we obtain asymptotically exact and consistent tests for general linear hypotheses in the parameters of the model. Additionally, we confirm that naive bootstrapping can not be used for covariance estimation in the pseudo-observation model either. However, it can be used for hypothesis testing by applying a suitable studentization. Simulations illustrate the good performance of our proposed methods in many scenarios. Finally, we obtain a general uniform law of large numbers for U- and V-statistics, as such statistics are central in the mathematical analysis of the inference procedures developed in this work.
@@ 1 Introduction
To investigate conditions under which the proposed method of Andersen et al. (2003) is valid, Graw et al. (2009) studied the pseudo-observations in the competing risks setting in terms of so called von-Mises expansions. However, it turned out that the remainder terms involved in this expansion cannot be ignored; while all other conclusions of Graw et al. (2009) remain valid, the asymptotic variance of the estimator contains second-order terms and is in general overestimated by the proposed Huber-White-type sandwich estimator of Andersen et al. (2003); see Jacobsen and Martinussen (2016) for details.
Overgaard et al. (2017) extended these results and obtained detailed conditions for the validity of the pseudo-observation approach, as well as the limiting variance of the estimators in the general case. They also conjectured that the usual sandwich estimator fails to correctly estimate the variance in general; instead, they suggested a plug-in estimator. Simulations suggest that this estimator performs well in many scenarios (Overgaard et al., 2018) but a mathematical proof of its consistency is not yet available. Another approach to variance estimation is the bootstrap; it offers another alternative to the Huber-White estimator. Although this was suggested more than twenty years ago (Andersen et al., 2003), the mathematical validity of such a procedure has still not been investigated in the present context either.
@@ 2 Model, notation, and previous results
for known matrices $C\in\mathbb{R}^{p\times q}$ and vectors $b\in\mathbb{R}^{p}$. Consider i.i.d. tuples $(V_{1},X_{1},Z_{1}),\dots,(V_{n},X_{n},Z_{n})$ which are copies of $(V,X,Z)$.
@@
Based on this representation, Overgaard et al. (2017) also established the identity $E(\hat{\theta}^{\ast}_{n,k}|Z_{k})=\phi(F)+E(\dot{\phi}(X_{k})|Z_{k})$. Therefore, a sufficient condition for the estimating equation (3) to be asymptotically unbiased, is $E(\hat{\theta}^{\ast}_{n,k}|Z_{k})=\mu(\beta_{0},Z_{k})$ which is implied by the assumption
@@
$E(\dot{\phi}(X)|Z)=E(V|Z)-\phi(F),$
if the mean function is correctly specified. For the remainder of this paper, we make the Assumptions 1 and 2. Based on these observations, Overgaard et al. (2017) obtained the asymptotic normality of
@@ 3 Covariance estimation and hypothesis tests
In general, $\Sigma$ and $\tilde{\Sigma}$ are not identical because $\tilde{\Sigma}$ ignores the second derivative $\ddot{\phi}$ induced by the essential part of the pseudo-observations. Furthermore, the diagonal elements of $\tilde{\Sigma}$ often exceed those of $\Sigma$ (Jacobsen and Martinussen, 2016; Overgaard et al., 2018; Overgaard, 2024). Therefore, the Huber-White estimator is biased. As the true variance is overestimated, hypothesis tests based on this estimator may lose a considerable amount of power. This is a particular problem in fields such as medicine where obtaining large samples can be challenging or expensive.
@@
These results demonstrate that hypothesis tests based on the Wald-type statistic can be applied in practice as long as the corrected variance estimator $\hat{\Sigma}_{n}^{PV}$ is available for the estimand of interest. From a theoretical perspective this is of no concern, but especially the explicit computation of the second-order influence function may be cumbersome in applications.
@@ 5.3 Simulation results
Table 2: Rejection rates under $H_{1}^{(2)}:C^{(2)}\beta_{0}\neq 0$, the true effect is $\delta_{2}=-1.$ The test with the highest power for each scenario is printed in bold.
@@ 6 Discussion and recommendation
In future research, we will also investigate other means of bootstrapping which have been suggested by Overgaard et al. (2018); Parner et al. (2023); this concerns randomly drawing with replacement from the original observations and then re-solving the estimating equations as well as combining this bootstrap with the pseudo-observations’ essential part representation.
@@ Appendix B Proofs and auxiliary mathematical results
by the law of large numbers for V-statistics, which is applicable because $A(\beta,Z)$ is locally dominated square integrable.
With a similar calculation as in the proof of Lemma 4, we can represent $\hat{\Sigma}^{(2)}_{ij}(\beta,F)$ with the function
@@ C.1 Further simulation results under the alternative hypothesis
Table 4: Rejection rates under $H_{1}^{(2)}:C^{(2)}\beta_{0}\neq 0$, the true effect is $\delta_{2}=1.$
@@
Table 5: Rejection rates under $H_{1}^{(1)}:C^{(1)}\beta_{0}\neq 0$, the true effect is $\delta_{1}=0.5.$
@@
Table 6: Rejection rates under $H_{1}^{(1)}:C^{(1)}\beta_{0}\neq 0$, the true effect is $\delta_{1}=1.$

### v2

@@ Abstract
The pseudo-observation method is regularly applied to time-to-event data. However, to date such analyses have relied on not formally verified statements or ad-hoc methods regarding covariance estimation. This paper strives to close this gap in the literature. To begin with, we demonstrate that the usual Huber-White estimator is not consistent for the limiting covariance of parameter estimates in pseudo-observation regression approaches. By confirming that a plug-in estimator can be used instead, we obtain asymptotically exact and consistent tests for general linear hypotheses in the parameters of the model. Additionally, we confirm that naive bootstrapping can not be used for covariance estimation in the pseudo-observation model either. However, it can be used for hypothesis testing by applying a suitable studentization. Simulations illustrate the good performance of our proposed methods in many scenarios. Finally, we obtain a general uniform law of large numbers for U- and V-statistics, as such statistics are central in the mathematical analysis of the inference procedures developed in this work.
@@ 1 Introduction
To investigate conditions under which the proposed method of Andersen et al. (2003) is valid, Graw et al. (2009) studied the pseudo-observations in the competing risks setting in terms of so called von-Mises expansions. However, it turned out that the remainder terms involved in this expansion cannot be ignored; while all other conclusions of Graw et al. (2009) remain valid, the asymptotic variance of the estimator contains second-order terms and is in general overestimated by the proposed Huber-White-type sandwich estimator of Andersen et al. (2003); see Jacobsen and Martinussen (2016) for details. Overgaard et al. (2017) extended these results and obtained detailed conditions for the validity of the pseudo-observation approach, as well as the limiting variance of the estimators in the general case. They also conjectured that the usual sandwich estimator fails to correctly estimate the variance in general; instead, they suggested a plug-in estimator. Simulations suggest that this estimator performs well in many scenarios (Overgaard et al., 2018) but a mathematical proof of its consistency is not yet available. Another approach to variance estimation is the bootstrap; it offers another alternative to the Huber-White estimator. Although this was suggested more than twenty years ago (Andersen et al., 2003), the mathematical validity of such a procedure has still not been investigated in the present context either.
@@ 2 Model, notation, and previous results
for known matrices $C\in\mathbb{R}^{p\times q}$ and vectors $b\in\mathbb{R}^{p}$. Consider independent and identically distributed (i.i.d.) tuples $(V_{1},X_{1},Z_{1}),\dots,(V_{n},X_{n},Z_{n})$ which are copies of $(V,X,Z)$.
@@
Based on this representation, Overgaard et al. (2017) also established the identity $E(\hat{\theta}^{\ast}_{n,k}|Z_{k})=\phi(F)+E(\dot{\phi}(X_{k})|Z_{k})$. Therefore, a sufficient condition for the estimating equation (3) to be asymptotically unbiased, is $E(\hat{\theta}^{\ast}_{n,k}|Z_{k})=\mu(\beta_{0},Z_{k})$ which is implied by the following assumption if the mean function is correctly specified.
@@
$E(\dot{\phi}(X)|Z)=E(V|Z)-\phi(F).$
For example, in the context of survival analysis Overgaard et al. (2019) demonstrated that this assumption is fulfilled in the case of conditionally independent censoring if certain estimators that have an inverse probability of censoring weighting representation are considered. To elaborate, the nonparametric estimator of $\theta$ is of the form
$\hat{\theta}_{n}=\frac{1}{n}\sum_{k=1}^{n}V_{k}\frac{\mathbbm{1}(C_{k}\geq\min(T_{k},t))}{\hat{G}_{n}(\min(T_{k},t)|Z_{k})}.$
Here, $(T_{1},C_{1},Z_{1}),\dots,(T_{n},C_{n},Z_{n})$ are i.i.d., $T$ non-negative event times, $C$ non-negative censoring times, $t$ a fixed time point of interest, $Z$ covariates influencing the censoring distribution, and $\hat{G}_{n}$ an estimator for the survival function $G$ of the censoring random variables $C$. The assumption of conditional independence translates to the independence of $T$ and $C$ conditional on the covariates $Z.$ A common choice for $\hat{G}_{n}$ is the Kaplan-Meier estimator (Kaplan and Meier, 1958) stratified by covariate levels if the assumption is made that the censoring distribution only depends on categorical covariates. No stratification corresponds to the assumption of completely independent censoring. Other common and suitable choices for the censoring distribution are the proportional hazards model (Cox, 1972) or the additive hazard regression model (Aalen, 1989) as long as the support of the covariates is suitably restricted; cf. Overgaard et al. (2019) for details.
For the remainder of this paper, we make the Assumptions 1 and 2. Based on these observations, Overgaard et al. (2017) obtained the asymptotic normality of
@@ 3 Covariance estimation and hypothesis tests
In general, $\Sigma$ and $\tilde{\Sigma}$ are not identical because $\tilde{\Sigma}$ ignores the second derivative $\ddot{\phi}$ induced by the essential part of the pseudo-observations. Furthermore, the diagonal elements of $\tilde{\Sigma}$ often exceed those of $\Sigma$ (Jacobsen and Martinussen, 2016; Overgaard et al., 2018; Overgaard, 2025). Therefore, the Huber-White estimator is biased. As the true variance is overestimated, hypothesis tests based on this estimator may lose a considerable amount of power. This is a particular problem in fields such as medicine where obtaining large samples can be challenging or expensive.
@@
These results demonstrate that hypothesis tests based on the Wald-type statistic can be applied in practice as long as the corrected variance estimator $\hat{\Sigma}_{n}^{PV}$ is available for the estimand of interest. From a theoretical perspective this is of no concern but especially the explicit computation of the second-order influence function may be cumbersome in applications. For example, we refer to Sections 3 and 4 in Overgaard et al. (2018) for the elaborate and explicit formulas of the theoretical covariances and the resulting plug-in estimators related to the Aalen-Johansen estimator in competing risks models. Such derivations might bear the risk of mathematical errors and flaws in software implementations when the involved functionals are very complex.
@@ 5.1 Data summary and model
We make the assumption of completely independent censoring, in which case the pseudo-observations can be computed based on the leave-one-out Kaplan-Meier estimators of the event time’s survival function.
@@ 5.3 Simulation results
Table 2: Power of the tests under $H_{1}^{(2)}:C^{(2)}\beta_{0}\neq 0$, the true effect is $\delta_{2}=-1.$ The test with the highest power for each scenario is printed in bold.
@@ 6 Discussion and recommendation
In future research, we will also investigate other means of bootstrapping which have been suggested by Overgaard et al. (2018); Parner et al. (2023); this concerns randomly drawing with replacement from the original observations and then re-solving the estimating equations as well as combining this bootstrap with infinitesimal jackknife pseudo-observations. Furthermore, we aim to provide ready-to-use software implementations for these approaches in the future.
@@ Appendix B Proofs and auxiliary mathematical results
by the law of large numbers for V-statistics, which is applicable because $A(\beta,Z)$ is locally dominated square integrable. With a similar calculation as in the proof of Lemma 4, we can represent $\hat{\Sigma}^{(2)}_{ij}(\beta,F)$ with the function
@@ C.1 Further simulation results under the alternative hypothesis
Table 4: Power of the tests under $H_{1}^{(2)}:C^{(2)}\beta_{0}\neq 0$, the true effect is $\delta_{2}=1.$
@@
Table 5: Power of the tests under $H_{1}^{(1)}:C^{(1)}\beta_{0}\neq 0$, the true effect is $\delta_{1}=0.5.$
@@
Table 6: Power of the tests under $H_{1}^{(1)}:C^{(1)}\beta_{0}\neq 0$, the true effect is $\delta_{1}=1.$